OpenStudy (anonymous):

Find the equation of a line that passes through the point (4,2) that is perpendicular to the line y = x. Show your work.

OpenStudy (anonymous):

\[y=4/3x\]

OpenStudy (anonymous):

can u help?

OpenStudy (koikkara):

Hello, Sweet @princess1298 I was wondering... how can i help you ?

OpenStudy (anonymous):

awww haha i need to find the equation but show my work how i got to the equation

OpenStudy (koikkara):

lol, how can i show your work, wen u did it ....alright, can u just post your answer here...so that i can sort it out ?

OpenStudy (anonymous):

hahaha well i dont know how to get an equation outta this!!!!!!

OpenStudy (koikkara):

Alright, have a look at.. http://www.mathsisfun.com/algebra/line-parallel-perpendicular.html, Does this help ?

OpenStudy (anonymous):

nothing shows up it says oops cant find the file

OpenStudy (koikkara):

ohh, means.... u can't access the links..., just a min...

OpenStudy (anonymous):

is that it?!!!!!!!!!!!!

OpenStudy (koikkara):

Not yet.... check out this example...

OpenStudy (koikkara):

@princess1298 So u have the equations, i guess so... !!

OpenStudy (anonymous):

so i have to put it in y=mx+b????

OpenStudy (koikkara):

Yeah, You are brave..

OpenStudy (anonymous):

hahaha y u say that! im stupid!!

OpenStudy (koikkara):

Hmm... Because, ref this... (attach files) says well that you are just so incredible girl/boy to sort it out...

OpenStudy (anonymous):

hahaha nah!!!!!!!!!!

OpenStudy (koikkara):

At the end what happened to me, you knoe, i just forgot what was the question...! So, how much you got \(y=??\)

OpenStudy (anonymous):

idk if y is 1 or 0

OpenStudy (koikkara):

lol, sweet princess, can u please try it in here step by step so i can find where or what u made wrong...? \(Cute~Your~smile~makes~you~look~!!\), Its not part of the question...lol

OpenStudy (anonymous):

lol one sec

OpenStudy (anonymous):

i got y=-3/4x=5 is that right?

OpenStudy (koikkara):

umm, \(Partially,~ it ~looks ~Correct~!!\)

OpenStudy (anonymous):

hahha ughhh!!!!!!!!!! where am i wrong!!!!!

OpenStudy (koikkara):

@Instagram Hey can u help her... because, i have a powercut in here just after 1 min... Thank you..!

OpenStudy (koikkara):

@princess1298 Like you, she too a brilliant, could help you out, when I am Out !!

OpenStudy (anonymous):

UMM OK!!!!!!!!!!! hahaha

OpenStudy (anonymous):

y=mx+b y=4x+2

OpenStudy (anonymous):

I'm not sure what we're doing here.

OpenStudy (anonymous):

finding the equation